OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (anonymous):

okay I understand that the formula is \[\cos \theta = u \times v \div \left| u \right|\left| v \right| \] I plug in everything and then finally get \[38\div \sqrt{13}\sqrt{80}\] Now I have no idea what to do after this...

OpenStudy (stamp):

@jazzie Read throuh this http://tutorial.math.lamar.edu/Classes/CalcII/DotProduct.aspx

OpenStudy (stamp):

u.v=38 |u||v|=sqrt(20*73) angle=arccos(38/sqrt(1460)) 6°

OpenStudy (anonymous):

Okay I understand now. Thank you !!